Expand the brackets and then simplify the following expression 3(2x + y) +6(5 - x)
1
Expert's answer
2020-06-07T14:15:40-0400

"3(2x + y) + 6(5 - x) = 3*2x + 3*y +6*5 - 6*x"


"= 6x + 3y + 30 - 6x" "= 3y + 30"